Assume that 'x' is a number chosen at random from the sets of integers between 1 and 15 inclusive, what is the probability that 'x' is a:
i) Single digit number
ii) Multiple of 5 or 7
iii) Complete square.

Answer by Mathtut(3670)

i) 9 possibles 1-9.... 9/15=3/5
:
ii)5,10,15 7,14....so 5 possibles 5/15=1/3
:
iii)if we are talking about a perfect square we would have 3 possibles 1,4,9...so 3/15=1/5